# revtex4-2Repair the float

### v4

@@ 1 Introduction
Figure 1: Trotterisation performance for the A2A-Ising model (with $j=256$; see Sec. 3 for definition), quantified in this figure with the average point-to-point absolute error in expectation values over $t=200\leavevmode\nobreak\ (2\pi g^{-1})$ (see Sec. 4.4 for definitions and more detailed error analyses; data from Fig. 7 (g)), breaks down rapidly after reaching a certain Trotter step size, observed as the large increase in the errors.
@@ 4.1.3 Stable Regions
Figure 4: Time-averages of dynamical signatures for each model compared across system size: (a–c) local observables and (d–f) simulation fidelity. Left to right, columns show results for A2A-Ising, Heisenberg, and Rabi-Dicke models, respectively, with line colours corresponding to system sizes given in legends in (a–c), and x-axes common to each column showing step sizes. $\langle.\rangle_{t}$ represents time averaging over periods $t=200\leavevmode\nobreak\ (2\pi g^{-1})$ (A2A-Ising), $t=50\leavevmode\nobreak\ (2\pi g^{-1})$ (Heisenberg), and $t=200\leavevmode\nobreak\ (2\pi g^{-1})$ (Rabi-Dicke). (a–c) Expectation values of example local observables: (a) magnetisation $\langle\braket{\bar{J}_{z}}\rangle_{t}$, (b) polarization of qubit one $\langle\braket{\sigma_{z}^{1}}\rangle_{t}$, and (c) normalised photon number $\langle\langle n\rangle\rangle_{t}/n_{j}$. Dotted silver lines are time averages of sampled ideal dynamics, sampled at intervals of $\tau$, to illustrate the emergence of sampling errors at large step sizes, even in the absence of digitisation errors. (d–f) While time-averaged simulation fidelities $\langle\mathcal{F}(\psi_{\text{dig}}(t),\psi_{\text{ide}}(t))\rangle_{t}$ do not show the same sudden threshold because of pre-threshold quasiperiodic oscillations, they show a secondary drop at the threshold. Where they occur (e.g., for large enough $j$ values in the A2A-Ising and Rabi-Dicke models), stable regions are readily observed in time-averaged signatures, and are generally observed at the same position independent of size.
@@ 4.2 System-Size Dependence
The first column of Fig. 4 shows the time averages of (a) $\braket{\bar{J}_{z}}$ values and (d) simulation fidelities for a selection of system sizes between $j=1$ and $j=256$ (labelled in Fig. 4 (a)), averaged over $t=200\leavevmode\nobreak\ (2\pi g^{-1})$. Thresholds in the expectation values appear around the same position for any system size considered, and at the same location as a secondary drop in the simulation fidelities which marks its transition from oscillatory to decaying behaviour. At the smallest sizes ($j\in\{1,2,3\}$), however, we observe neither the secondary drop in simulation fidelity, nor thresholds in the expectation values, although substantial smooth deviations from the ideal dynamics do start to appear at around the same position. In Appendix C.1, we plot the dynamics of $\braket{\bar{J}_{z}}$ expectation values for $j\in\{1,2,3,4,6,8\}$, showing that destruction of quasiperiodicity in the full dynamics only becomes clearly observable once $j\gtrsim 4$, where the sharper threshold also emerges in the time averages. The threshold then sharpens and converges in position as system size increases.
@@
The second column of Fig. 4 shows the time averages of (b) $\braket{\sigma_{z}^{1}}$ values and (e) simulation fidelities for the system sizes given in the Fig. 4 (b), and for $t=50\leavevmode\nobreak\ (2\pi g^{-1})$. The threshold position is again largely independent of system size, but identification of the threshold from these time-averaged quantities is more subtle, due to initial-state effects discussed in Sec. 4.1.1. In Appendix D.2, we provide time averaging and error analyses for various types of initial states, confirming that different initial state choices with varying amounts of pre-threshold Trotter errors still show similar changes in behaviour at a shared threshold. In addition, for the same system sizes considered here, our analyses of eigenvector statistics in Sec. 4.3 show that global signatures of the full Trotter step unitaries (that is, encompassing all initial states) also show the same distinctive threshold. Note that the two-qubit case has unit simulation fidelity (yellow line in Fig. 4 (e)), because it has no Trotter errors, even though time-averaged quantities still exhibit sampling errors at large steps.
@@
The third column of Fig. 4 shows the time averages of (c) scaled $\langle n\rangle/n_{j}$ values and (f) simulation fidelities for the system sizes labelled in Fig. 4 (c), and for $t=200\leavevmode\nobreak\ (2\pi g^{-1})$. Here, we observe a threshold in time-averaged dynamics even at the smallest size $j=0.5$, the Rabi model. In the Appendix E.1, we include the dynamical evolutions of various quantities for the Rabi model, which also already show dynamical signatures suggestive of quantum chaos, such as the destruction of quasiperiodicity.
@@ 4.3.1 Goodness-of-Fit Test for Eigenvector Statistics
Figure 5: Histograms showing eigenvector statistics for the $j=256$ A2A-Ising Trotter step unitary (TSU) for representative step sizes in different regimes: (a) $\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$ (before the threshold), (b) $\tau=0.4\leavevmode\nobreak\ (2\pi g^{-1})$ (on the threshold), (c) $\tau=0.5\leavevmode\nobreak\ (2\pi g^{-1})$ (quantum chaotic regime beyond the threshold), and (d) $\tau=0.7\leavevmode\nobreak\ (2\pi g^{-1})$ (stable region in the quantum chaotic regime). Agreement with random matrix theory (RMT) distributions, shown for different time-reversal symmetries by black dashed and dotted lines in each plot, provides strong evidence for quantum chaotic dynamics. However, while the visual agreement and disagreement in the quantum-chaotic and regular regimes, respectively, in (a) and (c) seem clear, apparently similar visual agreement in (d) would then be misleading, as it corresponds to a stable region where the behaviours of dynamical signatures clearly deviate from quantum chaotic, and even (b) shows reasonable visual agreement at a step size in the transition where the dynamics still exhibits clearly quasiperiodic behaviour. By contrast, the reduced chi-squared goodness-of-fit test statistic $X^{2}_{\rm RMT}$, compared here against the COE distribution and labelled in coloured text on each figure, provides quantitative and objective evidence for agreement or disagreement with RMT and successfully predicts the observation of quantum chaotic dynamics in each case: for example, $X^{2}_{\rm RMT}$ shows clear disagreement with quantum chaotic dynamics in both transition regions (b) and even stable regions (d).
@@ 4.3.2 Failure of visual comparisons versus $X^{2}_{\rm RMT}$
Figure 5 shows the eigenvector statistics of the A2A-Ising TSU calculated in the $J_{z}$ basis for $j=256$ and, respectively, for the step sizes: (a) $\tau=0.002\leavevmode\nobreak\ (2\pi g^{-1})$ before the threshold, (b) $\tau=0.4\leavevmode\nobreak\ (2\pi g^{-1})$ on the threshold, (c) $\tau=0.5\leavevmode\nobreak\ (2\pi g^{-1})$ in the quantum chaotic regime beyond the threshold, and (d) $\tau=0.7\leavevmode\nobreak\ (2\pi g^{-1})$ on a stable region in the quantum chaotic regime. Recall that collective spin behaviour in the A2A-Ising DQS restricts the system evolution into a total-spin conserved subspace $\mathcal{D}_{j}=2j+1$ of the full Hilbert space $\mathcal{D}_{N}=2^{N}$ of $N$ spin-1/2 systems. It therefore cannot display many-body quantum chaos of the full $N$-spin system, but can still exhibit the quantum chaos of a single spin $j$ within each symmetry subspace. The standard RMT distributions in Fig. 5 are therefore plotted for $\mathcal{D}_{j}=2j+1$. Because the A2A-Ising decomposition in Eq. (8) has non-symplectic time-reversal symmetry, it should show COE statistics [46]. While Figure 5 panels (a) and (c) show clear visual disagreement and agreement, respectively, with the expected COE distribution, the apparent visual agreement in Fig. 5 (d) is highly misleading, as this corresponds to a stable region, where the quantum chaotic dynamics are clearly interrupted. To characterise this agreement in a more quantitative manner, we introduce a reduced-chi-squared goodness-of-fit test statistic $X^{2}_{\rm RMT}$ to provide a generic quantitative signature for the agreement with the eigenvector statistics expected for random matrix theory (see Appendix A.3 for the definition and more details). The values of $X^{2}_{\rm RMT}$ comparing the histogram plots with the relevant COE distribution in Figs 5 (a–d) are, respectively, (a) $\sim 106084$, (b) $\sim 679$, (c) $\sim 1.001$, and (d) $\sim 333$, successfully identifying not only agreement and disagreement in the quantum-chaotic and regular regimes, but also clear disagreement with quantum chaotic eigenvector statistics within the stable region. Even Figure 5 (b) arguably also agrees reasonably well visually, but $X^{2}_{\rm RMT}$ clearly demonstrates there is still strong disagreement with the target COE distribution at this point in the middle of the transition.
@@
In general, RMT can be used to make specific quantitative predictions about the expected distributions of eigenvector coefficients for random unitary matrices, but does not provide any guidance as to the eigenvector distributions expected for non-random matrices. Similarly, the $X^{2}_{\rm RMT}$ goodness-of-fit test enables quantitatively rigorous analysis of whether the observed distribution for a particular unitary is consistent or inconsistent with the distribution predicted by RMT, with agreement requiring $X^{2}_{\rm RMT}\sim 1$. However, while $X^{2}_{\rm RMT}$ sufficiently greater than 1 can definitively rule out agreement with a particular model, the goodness-of-fit test makes no generic prediction about what $X^{2}_{\rm RMT}$ should be in the case of disagreement. In quantum chaotic systems, eigenvectors are generally quite evenly distributed across all basis states, typically giving $\eta\sim 1/\mathcal{D}\ll 1$). Except for very small systems, RMT probability densities rapidly approach exponentially small values for increasing $\eta$, with negligible support across most of $\eta$’s allowed range ($1/\mathcal{D}\ll\eta\leq 1$). Eigenvectors for regular system dynamics, however, commonly contain larger components ($\eta\gg 1/\mathcal{D}$) which can greatly distort $X^{2}$ goodness-of-fit comparisons against RMT distributions unless significant care is taken in the way the calculated values of $\eta$ are binned to form a histogram. At the same time, these are exactly the components we want to retain sensitivity to in order to detect deviations from quantum chaotic behaviour. In Appendix A.3, we provide a detailed initial recipe for how to implement the $X^{2}_{\rm RMT}$ goodness-of-fit test in a robust and rigorous way, which involves dividing up the main part of the comparison RMT probability distribution into equal probability bins ($\sim 10$ counts per bin) and capturing the exponentially vanishing tail region in a single, low-probability tail bin ($\ll 1$ count) that enhances sensitivity to small regions of stable dynamics in Hilbert space. This provides an important first step in the development of objective statistical techniques for identifying RMT, which we will study further in future work.
@@ A.4.1 Static Signatures
The eigenvector statistics are analysed using the prescription in Appendix A.3, noting the subtleties that arise when the system has degeneracies or additional physical symmetries. For example, when analysing eigenvector statistics for the CSE kicked top, which exhibits Kramers degeneracies, the additional eigenvector freedom introduced by these degeneracies can be taken into account by summing the eigenvector component populations for the degenerate eigenstates. This procedure decreases the number of components $\eta$ by half, but this is already taken into account in the derivation of the CSE distribution in Eq. (35[56]. If the system has an additional symmetry, such as the $R_{y}(\pi)$ rotational symmetries present in the COE and CUE kicked top models described above, the eigenvector statistics are analysed using only the eigenvector components within the symmetry subspaces. In such cases, the distributions in Eqs (3334) must be calculated using the decreased degree of freedom, here $\mathcal{D}/2$. The same approach is used to account for the parity symmetries present in the Heisenberg and Dicke model DQS systems studied in the rest of this paper.
@@ Appendix B Other Signatures of Quantum Chaos
Figure 10: Other dynamical signatures of quantum chaos (and their time-averages for each model compared across system size), showing (a–c) (and (g–i)) delocalisation and (d–f) (and (j–l)) further complementary signatures. Left to right (and for color plots), columns show results for A2A-Ising (with $j=64$), Heisenberg (chain of $N=8$ qubits), and Dicke (with $j=6$) models, respectively, with x-axes common to each column showing step sizes. The y-axes of (a–f) are the simulation times, and the specific quantities plotted on the y-axes of (g–l) are given in the sub-figure titles, with $\langle.\rangle_{t}$ representing time averaging over periods $t=200\leavevmode\nobreak\ (2\pi g^{-1})$ (A2A-Ising), $t=50\leavevmode\nobreak\ (2\pi g^{-1})$ (Heisenberg), and $t=200\leavevmode\nobreak\ (2\pi g^{-1})$ (Rabi-Dicke) and with line colours corresponding to system sizes given in legends in (g–i). The black lines (right y-axes) in (a–f) are the time averages of the colour plots. Dotted silver lines are time averages of sampled ideal dynamics, sampled at intervals of $\tau$, to illustrate the emergence of sampling effects at large step sizes, even in the absence of digitisation errors. In each model, dynamical quantities share a common Trotterisation threshold at a critical step size separating regular from quantum chaotic regimes. (a–c) Localised initial states rapidly delocalise beyond the threshold, quantified by participation ratios (PR, Eq. (43)). (d–f) further complementary dynamical signatures of quantum chaos exhibit the same threshold: e.g., infidelity under perturbation $1-\mathcal{F}(\psi_{p}(t),\psi(t))$ (see Eq. (44)) for Ising, and normalised sub-system entropies $\mathcal{S}(\rho_{i})$ (Eq. (45)) for Heisenberg (first qubit) and Dicke (spin), are all rapidly maximised beyond the threshold. (g–i) Time-averaged participation ratios $\langle\text{PR}\rangle_{t}$ show clear delocalisation beyond the threshold. (j–l) Beyond the threshold, infidelity under perturbation $\langle 1-\mathcal{F}(\psi_{p}(t),\psi(t))\rangle_{t}$ for Ising in (j), and normalised sub-system entropies $\langle\mathcal{S}(\rho_{i})\rangle_{t}$ for Heisenberg (first qubit) (k) and Dicke (spin) (l), rapidly approach maximum values for increasing system size. In each model, the dynamical signatures of quantum chaos (a–l) disappear in the stable regions, showing (at least partial) revival of regular dynamics.

### v5

# Quantum Chaos and Universal Trotterisation Behaviours in Digital Quantum Simulations

@@ 1 Introduction
Figure 1: Trotterisation performance for the A2A-Ising model (with $j=256$; see Sec. 3 for definition), quantified in this figure with the average point-to-point absolute error in expectation values over $t=200\penalty 10000\ (2\pi g^{-1})$ (see Sec. 4.4 for definitions and more detailed error analyses; data from Fig. 7 (g)), breaks down rapidly after reaching a certain Trotter step size, observed as the large increase in the errors.
@@ 4.1.3 Stable Regions
Figure 4: Time-averages of dynamical signatures for each model compared across system size: (a–c) local observables and (d–f) simulation fidelity. Left to right, columns show results for A2A-Ising, Heisenberg, and Rabi-Dicke models, respectively, with line colours corresponding to system sizes given in legends in (a–c), and x-axes common to each column showing step sizes. $\langle.\rangle_{t}$ represents time averaging over periods $t=200\penalty 10000\ (2\pi g^{-1})$ (A2A-Ising), $t=50\penalty 10000\ (2\pi g^{-1})$ (Heisenberg), and $t=200\penalty 10000\ (2\pi g^{-1})$ (Rabi-Dicke). (a–c) Expectation values of example local observables: (a) magnetisation $\langle\braket{\bar{J}_{z}}\rangle_{t}$, (b) polarization of qubit one $\langle\braket{\sigma_{z}^{1}}\rangle_{t}$, and (c) normalised photon number $\langle\langle n\rangle\rangle_{t}/n_{j}$. Dotted silver lines are time averages of sampled ideal dynamics, sampled at intervals of $\tau$, to illustrate the emergence of sampling errors at large step sizes, even in the absence of digitisation errors. (d–f) While time-averaged simulation fidelities $\langle\mathcal{F}(\psi_{\text{dig}}(t),\psi_{\text{ide}}(t))\rangle_{t}$ do not show the same sudden threshold because of pre-threshold quasiperiodic oscillations, they show a secondary drop at the threshold. Where they occur (e.g., for large enough $j$ values in the A2A-Ising and Rabi-Dicke models), stable regions are readily observed in time-averaged signatures, and are generally observed at the same position independent of size.
@@ 4.2 System-Size Dependence
The first column of Fig. 4 shows the time averages of (a) $\braket{\bar{J}_{z}}$ values and (d) simulation fidelities for a selection of system sizes between $j=1$ and $j=256$ (labelled in Fig. 4 (a)), averaged over $t=200\penalty 10000\ (2\pi g^{-1})$. Thresholds in the expectation values appear around the same position for any system size considered, and at the same location as a secondary drop in the simulation fidelities which marks its transition from oscillatory to decaying behaviour. At the smallest sizes ($j\in\{1,2,3\}$), however, we observe neither the secondary drop in simulation fidelity, nor thresholds in the expectation values, although substantial smooth deviations from the ideal dynamics do start to appear at around the same position. In Appendix C.1, we plot the dynamics of $\braket{\bar{J}_{z}}$ expectation values for $j\in\{1,2,3,4,6,8\}$, showing that destruction of quasiperiodicity in the full dynamics only becomes clearly observable once $j\gtrsim 4$, where the sharper threshold also emerges in the time averages. The threshold then sharpens and converges in position as system size increases.
@@
The second column of Fig. 4 shows the time averages of (b) $\braket{\sigma_{z}^{1}}$ values and (e) simulation fidelities for the system sizes given in the Fig. 4 (b), and for $t=50\penalty 10000\ (2\pi g^{-1})$. The threshold position is again largely independent of system size, but identification of the threshold from these time-averaged quantities is more subtle, due to initial-state effects discussed in Sec. 4.1.1. In Appendix D.2, we provide time averaging and error analyses for various types of initial states, confirming that different initial state choices with varying amounts of pre-threshold Trotter errors still show similar changes in behaviour at a shared threshold. In addition, for the same system sizes considered here, our analyses of eigenvector statistics in Sec. 4.3 show that global signatures of the full Trotter step unitaries (that is, encompassing all initial states) also show the same distinctive threshold. Note that the two-qubit case has unit simulation fidelity (yellow line in Fig. 4 (e)), because it has no Trotter errors, even though time-averaged quantities still exhibit sampling errors at large steps.
@@
The third column of Fig. 4 shows the time averages of (c) scaled $\langle n\rangle/n_{j}$ values and (f) simulation fidelities for the system sizes labelled in Fig. 4 (c), and for $t=200\penalty 10000\ (2\pi g^{-1})$. Here, we observe a threshold in time-averaged dynamics even at the smallest size $j=0.5$, the Rabi model. In the Appendix E.1, we include the dynamical evolutions of various quantities for the Rabi model, which also already show dynamical signatures suggestive of quantum chaos, such as the destruction of quasiperiodicity.
@@ 4.3.1 Goodness-of-Fit Test for Eigenvector Statistics
Figure 5: Histograms showing eigenvector statistics for the $j=256$ A2A-Ising Trotter step unitary (TSU) for representative step sizes in different regimes: (a) $\tau=0.002\penalty 10000\ (2\pi g^{-1})$ (before the threshold), (b) $\tau=0.4\penalty 10000\ (2\pi g^{-1})$ (on the threshold), (c) $\tau=0.5\penalty 10000\ (2\pi g^{-1})$ (quantum chaotic regime beyond the threshold), and (d) $\tau=0.7\penalty 10000\ (2\pi g^{-1})$ (stable region in the quantum chaotic regime). Agreement with random matrix theory (RMT) distributions, shown for different time-reversal symmetries by black dashed and dotted lines in each plot, provides strong evidence for quantum chaotic dynamics. However, while the visual agreement and disagreement in the quantum-chaotic and regular regimes, respectively, in (a) and (c) seem clear, apparently similar visual agreement in (d) would then be misleading, as it corresponds to a stable region where the behaviours of dynamical signatures clearly deviate from quantum chaotic, and even (b) shows reasonable visual agreement at a step size in the transition where the dynamics still exhibits clearly quasiperiodic behaviour. By contrast, the reduced chi-squared goodness-of-fit test statistic $X^{2}_{\rm RMT}$, compared here against the COE distribution and labelled in coloured text on each figure, provides quantitative and objective evidence for agreement or disagreement with RMT and successfully predicts the observation of quantum chaotic dynamics in each case: for example, $X^{2}_{\rm RMT}$ shows clear disagreement with quantum chaotic dynamics in both transition regions (b) and even stable regions (d).
@@ 4.3.2 Failure of visual comparisons versus $X^{2}_{\rm RMT}$
Figure 5 shows the eigenvector statistics of the A2A-Ising TSU calculated in the $J_{z}$ basis for $j=256$ and, respectively, for the step sizes: (a) $\tau=0.002\penalty 10000\ (2\pi g^{-1})$ before the threshold, (b) $\tau=0.4\penalty 10000\ (2\pi g^{-1})$ on the threshold, (c) $\tau=0.5\penalty 10000\ (2\pi g^{-1})$ in the quantum chaotic regime beyond the threshold, and (d) $\tau=0.7\penalty 10000\ (2\pi g^{-1})$ on a stable region in the quantum chaotic regime. Recall that collective spin behaviour in the A2A-Ising DQS restricts the system evolution into a total-spin conserved subspace $\mathcal{D}_{j}=2j+1$ of the full Hilbert space $\mathcal{D}_{N}=2^{N}$ of $N$ spin-1/2 systems. It therefore cannot display many-body quantum chaos of the full $N$-spin system, but can still exhibit the quantum chaos of a single spin $j$ within each symmetry subspace. The standard RMT distributions in Fig. 5 are therefore plotted for $\mathcal{D}_{j}=2j+1$. Because the A2A-Ising decomposition in Eq. (8) has non-symplectic time-reversal symmetry, it should show COE statistics [46]. While Figure 5 panels (a) and (c) show clear visual disagreement and agreement, respectively, with the expected COE distribution, the apparent visual agreement in Fig. 5 (d) is highly misleading, as this corresponds to a stable region, where the quantum chaotic dynamics are clearly interrupted. To characterise this agreement in a more quantitative manner, we introduce a reduced-chi-squared goodness-of-fit test statistic $X^{2}_{\rm RMT}$ to provide a generic quantitative signature for agreement with the expected RMT eigenvector statistics, with agreement requiring $X^{2}_{\rm RMT}\sim 1$ (see Appendix A.3 for the definition and more details). The level of agreement is quantified by comparing $X^{2}_{\rm RMT}$ with confidence regions for the reduced chi-squared distribution: these scale in size roughly inversely with the square root of the chi-squared degrees of freedom (here, $\text{DOFs}\sim\text{histogram bins}$, which we set proportional to $\mathcal{D}^{2}$). The values of $X^{2}_{\rm RMT}$ comparing the histogram plots with the relevant COE distribution in Figs 5 (a–d) are, respectively, (a) $\sim 106084$, (b) $\sim 679$, (c) $\sim 1.001$, and (d) $\sim 333$ (compared with a $95\%$ confidence region—approximately $\pm 2\sigma$—of 0.983–1.017 for $j=256$), successfully identifying not only agreement and disagreement in the quantum-chaotic and regular regimes, but also clear disagreement with quantum chaotic eigenvector statistics within the stable region. Even Figure 5 (b) arguably also agrees reasonably well visually, but $X^{2}_{\rm RMT}$ clearly demonstrates there is still strong disagreement with the target COE distribution at this point in the middle of the transition.
@@
In general, RMT can be used to make specific quantitative predictions about the expected distributions of eigenvector coefficients for random unitary matrices, but does not provide any guidance as to the eigenvector distributions expected for non-random matrices. Similarly, the $X^{2}_{\rm RMT}$ goodness-of-fit test enables quantitatively rigorous analysis of whether the observed distribution for a particular unitary is consistent or inconsistent with the distribution predicted by RMT. However, while $X^{2}_{\rm RMT}$ sufficiently greater than 1 can definitively rule out agreement with a particular model, the goodness-of-fit test makes no generic prediction about what $X^{2}_{\rm RMT}$ should be in the case of disagreement. In quantum chaotic systems, eigenvectors are generally quite evenly distributed across all basis states, typically giving $\eta\sim 1/\mathcal{D}\ll 1$). Except for very small systems, RMT probability densities rapidly approach exponentially small values for increasing $\eta$, with negligible support across most of $\eta$’s allowed range ($1/\mathcal{D}\ll\eta\leq 1$). Eigenvectors for regular system dynamics, however, commonly contain larger components ($\eta\gg 1/\mathcal{D}$) which can greatly distort $X^{2}$ goodness-of-fit comparisons against RMT distributions unless significant care is taken in the way the calculated values of $\eta$ are binned to form a histogram. At the same time, these are exactly the components we want to retain sensitivity to in order to detect deviations from quantum chaotic behaviour. In Appendix A.3, we provide a detailed initial recipe for how to implement the $X^{2}_{\rm RMT}$ goodness-of-fit test in a robust and rigorous way, which involves dividing up the main part of the comparison RMT probability distribution into equal probability bins ($\sim 10$ counts per bin) and capturing the exponentially vanishing tail region in a single, low-probability tail bin ($\ll 1$ count) that enhances sensitivity to small regions of stable dynamics in Hilbert space. This provides an important first step in the development of objective statistical techniques for identifying RMT, which we will study further in future work.
@@ A.4.1 Static Signatures
The eigenvector statistics are analysed using the prescription in Appendix A.3, noting the subtleties that arise when the system has degeneracies or additional physical symmetries. For example, when analysing eigenvector statistics for the CSE kicked top, which exhibits Kramers degeneracies, the additional eigenvector freedom introduced by these degeneracies can be taken into account by summing the eigenvector component populations for the degenerate eigenstates. This procedure decreases the number of components $\eta$ by half, but this is already taken into account in the derivation of the CSE distribution in Eq. (35) [56]. If the system has an additional symmetry, such as the $R_{y}(\pi)$ rotational symmetries present in the COE and CUE kicked top models described above, the eigenvector statistics are analysed using only the eigenvector components within the symmetry subspaces. In such cases, the distributions in Eqs (3334) must be calculated using the decreased degree of freedom, here $\mathcal{D}/2$. The same approach is used to account for the parity symmetries present in the Heisenberg and Dicke model DQS systems studied in the rest of this paper.
@@ Appendix B Other Signatures of Quantum Chaos
Figure 10: Other dynamical signatures of quantum chaos (and their time-averages for each model compared across system size), showing (a–c) (and (g–i)) delocalisation and (d–f) (and (j–l)) further complementary signatures. Left to right (and for color plots), columns show results for A2A-Ising (with $j=64$), Heisenberg (chain of $N=8$ qubits), and Dicke (with $j=6$) models, respectively, with x-axes common to each column showing step sizes. The y-axes of (a–f) are the simulation times, and the specific quantities plotted on the y-axes of (g–l) are given in the sub-figure titles, with $\langle.\rangle_{t}$ representing time averaging over periods $t=200\penalty 10000\ (2\pi g^{-1})$ (A2A-Ising), $t=50\penalty 10000\ (2\pi g^{-1})$ (Heisenberg), and $t=200\penalty 10000\ (2\pi g^{-1})$ (Rabi-Dicke) and with line colours corresponding to system sizes given in legends in (g–i). The black lines (right y-axes) in (a–f) are the time averages of the colour plots. Dotted silver lines are time averages of sampled ideal dynamics, sampled at intervals of $\tau$, to illustrate the emergence of sampling effects at large step sizes, even in the absence of digitisation errors. In each model, dynamical quantities share a common Trotterisation threshold at a critical step size separating regular from quantum chaotic regimes. (a–c) Localised initial states rapidly delocalise beyond the threshold, quantified by participation ratios (PR, Eq. (43)). (d–f) further complementary dynamical signatures of quantum chaos exhibit the same threshold: e.g., infidelity under perturbation $1-\mathcal{F}(\psi_{p}(t),\psi(t))$ (see Eq. (44)) for Ising, and normalised sub-system entropies $\mathcal{S}(\rho_{i})$ (Eq. (45)) for Heisenberg (first qubit) and Dicke (spin), are all rapidly maximised beyond the threshold. (g–i) Time-averaged participation ratios $\langle\text{PR}\rangle_{t}$ show clear delocalisation beyond the threshold. (j–l) Beyond the threshold, infidelity under perturbation $\langle 1-\mathcal{F}(\psi_{p}(t),\psi(t))\rangle_{t}$ for Ising in (j), and normalised sub-system entropies $\langle\mathcal{S}(\rho_{i})\rangle_{t}$ for Heisenberg (first qubit) (k) and Dicke (spin) (l), rapidly approach maximum values for increasing system size. In each model, the dynamical signatures of quantum chaos (a–l) disappear in the stable regions, showing (at least partial) revival of regular dynamics.